# Lost in Cultural Translation: Do LLMs Struggle with Math Across Cultural Contexts?

### v1

@@ 5 Prompt
Our prompt is straightforward—it does not instruct the LLM to solve math problems in any specific way, nor does it emphasize math, numerical reasoning, or cultural elements. Instead, it simply guides the LLM to solve the problem step by step. The actual math problem is then introduced within this structured approach. We also use the same prompt on all of the models for all the datasets.
@@ 7 Performance Gap
The Table A3 4 explains the difference in performance (accuracy) of all the 14 models across all the datasets. These gaps are explained in detail in results section. The values in the table are presented in decimal form, whereas in the main paper (Graph 5), they have been converted to percentages for ease of interpretation.
@@ 8 Difference in Tokenization
We utilize the OpenAI tokenizer to analyze how a question from the original GSM8K dataset is tokenized compared to its Moldovan-contextualized version. Notably, the only modification in this example is the change of names, yet the tokenization process treats them differently. This discrepancy highlights how even minor cultural adaptations can alter tokenization, leading to variations in how the model processes and understands the question.
@@ 9 McNemar Test
  • A low p-value (p $<$ 0.05) means the accuracy drop is statistically significant and not due to random chance.
@@
  • LLaMA 3.1-70B consistently had high b-values across all datasets, meaning it frequently failed on culturally adapted versions of the questions while performing well on the original GSM8K. It had some of the lowest p-values (p $<$ 0.01), particularly in Moldova, Pakistan, Somalia, and Suriname. This suggests that the model’s pretraining data might not be diverse enough to handle different cultural contexts in mathematical reasoning.

### v2

@@ 6 Prompt
Our prompt is straightforward, it does not instruct the LLM to solve math problems in any specific way, nor does it emphasize math, numerical reasoning, or cultural elements. Instead, it simply guides the LLM to solve the problem step by step. The actual math problem is then introduced within this structured approach. We also use the same prompt on all of the models for all the datasets.
@@ 8 Performance Gap
The Table A4 explains the difference in performance (accuracy) of all the 14 models across all the datasets. These gaps are explained in detail in Section  5.1. The values in the table are presented in decimal form, whereas in the main paper Figure 4, they have been converted to percentages for ease of interpretation.
@@ 9 Difference in Tokenization
We utilize the OpenAI tokenizer to analyze how a question from the original GSM8K dataset is tokenized compared to its Moldovan-contextualized version. Notably, the only modification in this example is the change of names, yet the tokenization process treats them differently. This discrepancy highlights how even minor cultural adaptations can alter tokenization, suggesting that culturally adapted vocabulary may be processed differently by the model, though the direct impact on reasoning performance cannot be established from tokenization differences alone.
@@ 10 McNemar Test
  • A low p-value (p << 0.05) means the accuracy drop is statistically significant and not due to random chance.
@@
  • LLaMA 3.1-70B consistently had high b-values across all datasets, meaning it frequently failed on culturally adapted versions of the questions while performing well on the original GSM8K. It had some of the lowest p-values (p << 0.01), particularly in Moldova, Pakistan, Somalia, and Suriname. This suggests that the model’s pretraining data might not be diverse enough to handle different cultural contexts in mathematical reasoning.
@@ 11.3 Problem Interpretation Errors
11.4 Mathematical Reasoning Errors
Across our evaluation, we observed mathematical reasoning errors where cultural adaptation caused models to fundamentally alter their solution strategy despite the underlying logic remaining unchanged. A representative example involves a gift bag planning problem where the 0.75 bags-per-guest ratio already accounts for the 1/4 of guests who do not show up. In the original GSM8K version, the model correctly applies this ratio directly to the 16 invited guests, computing 0.75 × 16 = 12 bags and arriving at a total cost of \$24. However, in the Surinamese adaptation, the model introduces a redundant step, first computing expected attendees by applying the 1/4 no-show rate to arrive at 12, and then applying the 0.75 ratio again to this already-reduced figure, producing 9 bags and a total cost of SRD 18. The same mathematical structure that was handled correctly in the original framing breaks down once the cultural context shifts.
Figure A9: Claud-3.5-sonnet Reasoning
11.5 Calculation Errors
Unlike reasoning errors where the solution strategy itself breaks down, the calculation errors we observed involve models that correctly structured their approach but produced wrong final answers due to arithmetic mistakes introduced after cultural adaptation. A representative example involves a week-long attendance counting problem where the original GSM8K version is solved correctly, with the model accurately summing 50, 70, 140, and 240 people across the seven days to arrive at 500. In the Somalian adaptation, replacing the character name and substituting “beach” with “Tuulo”, the model correctly computes the count for each individual day yet arrives at an incorrect total of 400, omitting the day one count from the final summation. The same arithmetic, performed correctly in the original framing, breaks down once the cultural context shifts.
Figure A10: Qwen 2.5-32B-Instruct reasoning
11.6 Other / unclear errors
A small set of errors in our analysis cannot be attributed to mathematical reasoning or calculation failures, and instead arise from issues such as answer extraction mismatches, formatting inconsistencies, and non-numeric outputs. For example, in a Moldovan-adapted history test scoring problem, the model correctly calculates 8 points from multiple-choice questions, 18 from true/false questions, and 15 from long-answer questions, explicitly stating a correct total of 41 points within its reasoning trace, yet the final answer field returns 3, the number of correctly answered long-answer questions rather than the total score.
Figure A11: Claud-3.5-sonnet Reasoning
12 Quantitative Error Analysis
To systematically quantify error patterns across all models and cultural variants, we conducted a comprehensive annotation of all incorrect model responses identified in our evaluation. We collected all cases where at least one of the three model runs produced an incorrect answer according to the strict consistency criterion described in Section  5.1. This resulted in 18,887 error instances across 14 models and 7 datasets (original GSM8K plus 6 cultural variants).
We removed cases involving only minor rounding or formatting differences that did not reflect genuine reasoning errors. Duplicate entries for the same model-culture-question combination were also removed to ensure each unique error was counted only once. Each error was categorized into one of six types: (1) Calculation error, (2) Mathematical Reasoning error, (3) Relationship misunderstanding, (4) Unit/currency error, (5) Problem misinterpretation, and (6) Other/unclear. The categorization was performed using an LLM-assisted annotation script that analyzed the model’s reasoning process for the first failing run among the three attempts.
To ensure annotation quality, we manually reviewed 50 randomly sampled error cases. The manual review confirmed that the LLM-based categorization was largely accurate, with minor discrepancies primarily occurring in the ”Other/unclear” category, which represents only 0.76% of total errors.
12.1 Error Distribution by Culture
Error Category GSM8K Hti Mld Pak Sol Som Sur
Mathematiical Reasoning error 1362 1470 1475 1490 1525 1508 1501
(58.58%) (54.63%) (54.61%) (53.20%) (54.78%) (53.49%) (54.27%)
Calculation error 890 918 893 1002 881 1007 924
(38.62%) (34.11%) (33.06%) (35.77%) (31.65%) (35.72%) (33.41%)
Relationship misunderstanding 65 119 140 103 195 85 144
(2.80%) (4.42%) (5.18%) (3.68%) (7.00%) (3.02%) (5.21%)
Unit / currency error 0 137 129 113 97 146 124
(0%) (5.09%) (4.78%) (4.03%) (3.48%) (5.18%) (4.48%)
Problem misinterpretation 0 23 47 65 66 40 51
(0%) (0.85%) (1.74%) (2.32%) (2.37%) (1.42%) (1.84%)
Other / unclear 0 24 17 28 20 33 22
(0%) (0.89%) (0.63%) (1.00%) (0.72%) (1.17%) (0.80%)
Table A7: Error distribution by cultural variant. Values show count with percentage in parentheses. Hti = HaiGSM8K, Mld = MolGSM8K, Pak = PakGSM8K, Sol = SolIGSM8K, Som = SomGSM8K, Sur = SurGSM8K.
Table  A7 reveals notable variation in culturally-specific error patterns across datasets. Solomon Islands exhibits the highest proportion of relationship misunderstandings (7.0%), consistent with the use of unfamiliar kinship terms such as ”tambu man.” Somalia and Haiti show elevated unit/currency errors (5.18% and 5.09%, respectively), reflecting challenges with less common currency units like the Somali Shilling and Haitian Gourde. The original GSM8K dataset shows the lowest rates of these culturally-specific errors, as expected.
12.2 Error Distribution by Model
Model Calculation Reason Relation Currency Misinterp Other/Unclear
Claude 3.5 185 163 13 36 13 4
(44.69%) (39.37%) (3.14%) (8.70%) (3.14%) (0.97%)
DeepSeek 285 305 55 52 20 9
(39.26%) (42.01%) (7.58%) (7.16%) (2.75%) (1.24%)
Gem Flash 2.0 214 325 37 23 20 5
(34.29%) (52.08%) (5.93%) (3.69%) (3.21%) (0.80%)
Gem Flash 1.5-8B 507 883 54 61 22 8
(33.03%) (57.52%) (3.52%) (3.97%) (1.43%) (0.52%)
Gemma 2-27B 538 620 29 50 20 10
(42.46%) (48.93%) (2.29%) (3.95%) (1.58%) (0.79%)
Gemma 2-9B 686 775 51 56 17 3
(43.20%) (48.80%) (3.21%) (3.53%) (1.07%) (0.19%)
LLaMA 3.1-70B 848 2211 117 68 20 12
(25.89%) (67.49%) (3.57%) (2.08%) (0.61%) (0.37%)
LLaMA 3.1-8B 840 2217 110 68 23 14
(25.67%) (67.76%) (3.36%) (2.08%) (0.70%) (0.43%)
Phi-3 Medium 780 1040 90 63 37 39
(38.07%) (50.76%) (4.39%) (3.07%) (1.81%) (1.90%)
Phi-4 345 336 84 54 22 3
(40.88%) (39.81%) (9.95%) (6.40%) (2.61%) (0.36%)
Mistral Large 299 365 52 50 20 19
(37.14%) (45.34%) (6.46%) (6.21%) (2.48%) (2.36%)
Mistral Saba 436 465 65 59 15 2
(41.84%) (44.63%) (6.24%) (5.66%) (1.44%) (0.19%)
GPT-4o 266 214 43 47 19 7
(44.63%) (35.91%) (7.21%) (7.89%) (3.19%) (1.17%)
Qwen 2.5-32B 294 412 51 59 24 9
(34.63%) (48.53%) (6.01%) (6.95%) (2.83%) (1.06%)
Table A8: Error distribution by model. Reason = Mathematical reasoning error, Relation = Relationship misunderstanding, Currency = Unit/currency error, Misinterp = Problem misinterpretation.
Table  A8 shows that error distribution patterns vary across models. Claude 3.5 Sonnet (C3.5) exhibits a notably higher proportion of calculation errors (44.69%) compared to reasoning errors (39.37%), contrasting with most other models where reasoning errors dominate. Claude 3.5 also shows the highest rate of unit/currency errors (8.70%), suggesting sensitivity to unfamiliar numerical contexts despite overall strong performance. Smaller models like LLaMA 3.1-8B show the highest absolute error counts, consistent with their lower overall accuracy reported in Section 4.2.2. Across all models, reasoning and calculation errors account for over 80% of failures, with culturally-specific errors (relationship and currency) comprising a smaller but consistent proportion.
12.3 Error Distribution by Entity Category
To complement the error type analysis in Section 12.1 and 12.2, Table A9 examines which cultural entity categories appear disproportionately more frequently in incorrect responses relative to their overall occurrence across the 1,198 culturally adaptable questions.
Category Occurrence % Error % Difference
Commerce and Economy 35.2% 44.8% +9.6%
Food and Drink 14.4% 22.2% +7.8%
Other 3.6% 4.1% +0.5%
Education 0.9% 1.2% +0.3%
Places and Locations 10.7% 10.9% +0.2%
Clothing and Appearance 1.6% 1.6% 0.0%
People and Roles 90.7% 89.2% -1.5%
Culture and Arts 5.6% 4.1% -1.5%
Table A9: Error Distribution by Cultural Entity Category
The Difference column indicates whether a category appears disproportionately more or less in incorrect responses relative to its overall occurrence. A positive value indicates the category is associated with more errors than its frequency alone would predict. Commerce and Economy entities show the largest positive difference (++9.6%), followed by Food and Drink (++7.8%), consistent with the currency-related and cultural entity interpretation errors, identified in the qualitative analysis in Section 5.2.4. In contrast, People and Roles shows a slightly negative difference (-1.5%), indicating that despite person names appearing in over 90% of questions, their substitution does not disproportionately cause errors. Note that this category is dominated by Person name entities (87.1% of questions). This dominance dilutes the signal from family member entities (4.9%) that were identified as error-prone in the qualitative analysis.